Magic Realm, Inc., has developed a new fantasy board game. The company sold 28,500 games last year at a selling price of $66 per game. Fixed expenses associated with the game total $475,000 per year, and variable expenses are $46 per game. Production of the game is entrusted to a printing contractor. Variable expenses consist mostly of payments to this contractor. Required: 1-a. Prepare a contribution format income statement for the game last year. 1-b. Compute the degree of operating leverage. 2. Management is confident that the company can sell 36,195 games next year (an increase of 7,695 games, or 27%, over last year). Given this assumption: a. What is the expected percentage increase in net operating income for next year? b. What is the expected amount of net operating income for next year? (Do not prepare an income statement; use the degree of operating leverage to compute your answer.)
Answer #1

a.

Contribution Format Income Statement
Total Cost Per Unit
Sales 1,881,000 66
Variable cost -1,311,000 -46
Contribution Margin 570,000 20
Fixed cost -475,000
Net Operating Income 95,000

b.

Degree of operating leverage = Contribution Margin / Net operating income

= 570,000/95,000

= 6

2.

Increase in sales = 27%

Increase in net operating income = Degree of operating leverage x Increase in sales

= 6 x 27

= 162%

2 b.

Expected net operating income for next year = 95,000 + 95,000 x 162%

= 95,000+153,900

= $248,900

Kindly comment if you need further assistance. Thanks‼!